OpenStudy (anonymous):

the total number of positive integral solutions of the equation x1.x2.x3.x4=10!

OpenStudy (anonymous):

yes

OpenStudy (perl):

x1.x2.x3.x4=3628800

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix @Compassionate @iambatman

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

no one knows the answer i think

OpenStudy (anonymous):

@Jhannybean

OpenStudy (arabpride):

Nope -No idea tbh~ cx

OpenStudy (jhannybean):

Is the way @perl solved it wrong?..

OpenStudy (anonymous):

yes it is wrong

ganeshie8 (ganeshie8):

As a start : find the prime factorization of 10!

OpenStudy (anonymous):

then after that

ganeshie8 (ganeshie8):

It will be easy if I walk you through

ganeshie8 (ganeshie8):

whats the prime factorization of 10! ?

ganeshie8 (ganeshie8):

thats not prime factorization

OpenStudy (anonymous):

its factorial

ganeshie8 (ganeshie8):

\[\rm 10! = 2^8\times 3^4\times 5^2\times 7^1\]

OpenStudy (anonymous):

yes then

ganeshie8 (ganeshie8):

Consider a divisor of \(10!\), It will be of form : \[\rm 2^a\times 3^b\times 5^c\times 7^d\] where \(\rm 0\le a\le 8, ~0\le b\le 4, ~0\le c\le 3, ~0\le d\le 1\) yes ?

OpenStudy (anonymous):

ok.then

OpenStudy (anonymous):

o<=c<=2

ganeshie8 (ganeshie8):

Oh rihgt, thanks for correcting @Princer_Jones :)

OpenStudy (anonymous):

ok.then

ganeshie8 (ganeshie8):

we're not done yet, we need to count number of solutions for (a,b,c,d) based on these divisors

OpenStudy (anonymous):

@ganeshie8 I am learning...:)

OpenStudy (anonymous):

its good princer

OpenStudy (anonymous):

ganeshie i m waiting

OpenStudy (jhannybean):

Patience, young grasshopper.

OpenStudy (anonymous):

ok miss jhanny

OpenStudy (anonymous):

ok.every question i post here is tricky bt i hope u will solve it

OpenStudy (anonymous):

jhanny u cn also try

ganeshie8 (ganeshie8):

let me tag few others @nerdguy2535 @ChristopherToni @dan815

OpenStudy (anonymous):

@rishavraj

ganeshie8 (ganeshie8):

Say \[\rm x_i = 2^{a_i}\times 3^{b_i}\times 5^{c_i}\times 7^{d_i} \] and look at your equation \[\large \rm x_1\cdot x_2\cdot x_3\cdot x_4 = 2^8\cdot 3^4\cdot 5^2\cdot 7^1\] Clearly each exponent of a prime, \(p\) in product \(\rm x_1\cdot x_2\cdot x_3\cdot x_4\) must equal to the exponent of \(p\) in \(10!\) : \[a_1 + a_2 + a_3 + a_4 = 8 \\ b_1 + b_2 + b_3 + b_4 = 4 \\ c_1+c_2+c_3+c_4 = 2 \\ d_1+d_2+d_3+d_4 = 1 \] find the number of nonnegative solutions of each of above equations and multiply them.

ganeshie8 (ganeshie8):

using stars and bars you get \[\large \rm \binom{11}{3}\cdot \binom{7}{3}\cdot \binom{5}{3}\cdot \binom{4}{3}\]

OpenStudy (dan815):

umm prolly too late, but just to confirm, you take 10! take all its prime factors, and see how many ways to get unique arrangements of that, then take stars and bars for each arrangement - then subtract the possible overlaps just wondering if there is a way in this too, i feel like the overlaps get ugly, i do like this solution though its nice say u have 2,2,2,3,3,3,5,5,7,7,7,7 _ _ _ _ . and you throw the number 2s around in there. For each arrangement of 2s in the bars there a different arrangemnt of 3s and 5s and 7s that will all multiply to 10! so the answer is (number of ways to split up the 3s between 4 bars) * (#of2s btwn 4 bars) * (#7btwen4bars...* and so on)

OpenStudy (anonymous):

@ganeshie8 can you explain the method, or post me any link of the theory section concerning this kind of problems.?? Thanks.:)

OpenStudy (dan815):

the 2nd part of my post is what ganeshie did

OpenStudy (dan815):

the process of splitting the number of 2s between 4 places is, the stars and bars method, like u have 8 2s, so p1+p2+p3+p4 = 8, where p1 to p4 cannot have a negative integer

OpenStudy (dan815):

2 2 2 2 2 2 2 2 <------------8 2s how to split into 4 place 2 2 2 | 2|2 2 | 2 2 <--------introduce these 3 bars, or 3 extra places so 11 Choose 3, u think of it as placing the 3 bars around in any of the 11 places now as there are 11 objects with the 2s and the bars in total now

OpenStudy (dan815):

so thats the total number of ways to split up the 2s, and for each split of 2 there are a differnt number of splits of 3s 5s and 7s that work, so u are multipling those arrangements together with this to get the total number of possible ways

ganeshie8 (ganeshie8):

@Princer_Jones i like this explanation here about stars and bars https://brilliant.org/discussions/thread/stars-and-bars/

ganeshie8 (ganeshie8):

lets do an exaple maybe

ganeshie8 (ganeshie8):

find the number of nonnegative integral solutions to the equation : \[\large \rm x_1 + x_2 + x_3 = 4\]

OpenStudy (anonymous):

OK...

ganeshie8 (ganeshie8):

we translate this problem by creating a bijection to a similar "stars and bars" problem

OpenStudy (anonymous):

there were sequences of length 13 in that problem

ganeshie8 (ganeshie8):

yes lets work the same problem maybe \[\large \rm a+b+c+d = 10\]

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

finding number of nonnegative integral solutions to above equation is equivalent to finding number of ways of arranging bars : \[\large \rm ***|***|***|*\]

ganeshie8 (ganeshie8):

number of stars before first bar gives you the value of \(a\), number of stars before first bar and second bar gives you the value of \(b\) etc..

ganeshie8 (ganeshie8):

In that string of stars and bars of length 13, you can choose 3 positions for bars in \(\large \binom{13}{3}\) ways.

OpenStudy (anonymous):

***|***|***|* are we to put the bars in this fashion only?

ganeshie8 (ganeshie8):

that gives one solution : (a,b,c,d) = (3,3,3,1)

ganeshie8 (ganeshie8):

\[\large \rm **|****|*|*** ~~~~\sim~~~ (2,4,1,3) \]

ganeshie8 (ganeshie8):

\[\large \rm **|****|****| ~~~~\sim~~~ (2,4,4,0) \]

OpenStudy (anonymous):

oh...

ganeshie8 (ganeshie8):

there is an one-to-one association between "stars and bars" problem and the equation we are solving

OpenStudy (anonymous):

hmm correct..

ganeshie8 (ganeshie8):

there are 10 stars and 3 bars, so the length of string is : 10+3 = 13

ganeshie8 (ganeshie8):

you can choose any 3 locations from those 13 locations for placing ur bars

OpenStudy (anonymous):

hmmm next?

ganeshie8 (ganeshie8):

the total number of different strings with 10 stars and 3 bars is : \[\large \rm \binom{13}{3}\] we're done.

OpenStudy (anonymous):

oh... this kind of problem falls under which branch of mathematics?

ganeshie8 (ganeshie8):

number theory

OpenStudy (dan815):

under life

ganeshie8 (ganeshie8):

basically what we're doing is translating an "unknown problem"(# of solutions to equation) to a known problem("stars and bars") by creating a bijection between the two

OpenStudy (anonymous):

oh i see.. :)....

ganeshie8 (ganeshie8):

lets do another example if u have time

OpenStudy (anonymous):

yeah i hv time..

ganeshie8 (ganeshie8):

\[\large \rm x_1 + x_2 + x_3 = 4\]

ganeshie8 (ganeshie8):

find number of nonnegative integral solutions to above eqn

ganeshie8 (ganeshie8):

now that we have "stars and bars" trick, we just translate the problem and follow our noses

ganeshie8 (ganeshie8):

how many stars ? how many bars ?

OpenStudy (dan815):

find number of nonnegative integer* solutions to above eqn

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4stars and 2 bars i gues..

ganeshie8 (ganeshie8):

yes because you want the left hand side things to add up to 4 you need bars to divide the stars

ganeshie8 (ganeshie8):

yes! \[\large \rm *|**|*\]

OpenStudy (anonymous):

hmm It should be 6C2 ???

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

you can choose 2 positions for our bars from the available 6 positions in 6C2 ways

OpenStudy (anonymous):

first one 13C2 and second one 9C2/@dan815

OpenStudy (dan815):

no the point is the bar is cutting both sets of stars at hte same time

OpenStudy (anonymous):

@dan815 i cant see the picture anymore.. and couldnot get your point..

OpenStudy (dan815):

-.-

OpenStudy (anonymous):

@ganeshie8 so using this stars and bars method we find the number of possible combinations of the 4 linear equations you establshed before, and they multiply them to get the final solution?

ganeshie8 (ganeshie8):

me neither

OpenStudy (dan815):

cuz im making stuff up

ganeshie8 (ganeshie8):

yes, you will feel "stars and bars" pretty mechanical once you get use to it...

ganeshie8 (ganeshie8):

\[a_1 + a_2 + a_3 + a_4 = 8 \\ b_1 + b_2 + b_3 + b_4 = 4 \\ c_1+c_2+c_3+c_4 = 2 \\ d_1+d_2+d_3+d_4 = 1\]

ganeshie8 (ganeshie8):

I bet you can work all above equations without much thinking

OpenStudy (anonymous):

Yeah obviously.:)

OpenStudy (anonymous):

(8+3 C 3)=11C3, (4+3C3)=7C3,...

ganeshie8 (ganeshie8):

\[\large \rm \binom{11}{3}\cdot \binom{7}{3}\cdot \binom{5}{3}\cdot \binom{4}{3}\]